OpenStudy (anonymous):

simplify by factoring sqrt 216x ^4

OpenStudy (saifoo.khan):

\[\Large x^2 6 \sqrt 6\]

OpenStudy (anonymous):

i have some others would you be able to help?

OpenStudy (saifoo.khan):

Sure, post it. How many do u have?

OpenStudy (anonymous):

just a few i think that i dont get 38 all together but i got most of them

OpenStudy (saifoo.khan):

Ok, post them here.. i will help u as much i can.

OpenStudy (anonymous):

okay this one i dont get there is a small 4 before the sqrt thing so i guess that is cube then sqrt 324x^8y^10

OpenStudy (saifoo.khan):

If u want u can post a screen shot here as well to avoid any mistakes.

OpenStudy (saifoo.khan):

\[\Large \sqrt[4]{324x^8y^{10}}\]That?

OpenStudy (anonymous):

yeah

OpenStudy (saifoo.khan):

\[\Large 3 x^2y^2 \sqrt[4]{2 y^2}\]

OpenStudy (anonymous):

ok give me a sec for the next one

OpenStudy (anonymous):

^4sqrt80x^12y^14

OpenStudy (anonymous):

written like the last one

OpenStudy (saifoo.khan):

\[\Large 3x^3y^3 \sqrt[4]{y^2}\]

OpenStudy (anonymous):

^5 sqrt 160x ^12y ^20

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

some of them i was able to do a few of them have me so lost

OpenStudy (anonymous):

that last one is telling me to factor into perfect fourths when i typed in your answer

OpenStudy (saifoo.khan):

\[\Large 2x^2y^4 \sqrt[5]{5x^2}\]

OpenStudy (saifoo.khan):

Nice, Mathlab is easy as cake. ;) A person can score As easily in ALgebra.

OpenStudy (saifoo.khan):

Which one? are u typing it in correct format?

OpenStudy (saifoo.khan):

Next question?

OpenStudy (anonymous):

i dont know how the one where you came up with 3x^3y^3 the little 4 in the sqrt y^2